# TrojFST: Embedding Trojans in Few-shot Prompt Tuning

### 25 Jan 2024 v2

@@ 6 Potential Defense
Bert-Large $74.53$ $71.38$ $98.78$ $42.73$
RoBERTa-Large $77.47$ $73.68$ $99.27$ $40.92$
Google T5-Base $79.70$ $75.94$ $99.63$ $53.08$
To the best of our knowledge, there are only few studies that specifically address defenses against backdoor attacks in NLP. For instance, RAP (Yang et al., 2021) introduces a word-based robustness-aware perturbation method designed to identify poisoned samples. And ONION (Qi et al., 2020) attempts to remove trigger words by empirically assessing sentence perplexities. However, they cannot handle our TrojFST using invisible syntactic triggers.
We propose a potential defense technique against TrojFST that minimizes its ASR by selectively pruning unimportant prompt tokens, assuming the defender knows the potential presence of a backdoored prompt. However, the token’s importance may vary with different input samples. Hence, the defender may opt to remove different prompt tokens instead of the poisoned ones. Therefore, even after token pruning, TrojFST can still achieve an ASR of over $40\%$, as demonstrated in Table 10. There is a need for a more efficient and accurate defense method.
@@ 7 Conclusion
In this paper, we propose a prompt-based backdoor attack, TrojFST, against PLMs through few-shot prompt-tuning. Instead of a full-model fine-tuning, TrojFST freezes the PLM, and trains a backdoored prompt for the PLM with only few input samples by tuning only one prompt token. The PLM remains clean throughout the entire TrojFST attack, making TrojFST more stealthy and resistant to existing encoder backdoor detection techniques. We also discuss the potential defense techniques in this paper. Compared to prior prompt-based backdoor attacks, TrojFST improves the ASR by $9\%\sim 48\%$ and the CDA by $4\%\sim 9\%$ across various PLMs and a wide range of downstream tasks.

### 19 Mar 2024 v3

@@ 5.2 Potential Defense
RoBERTa-Large $77.5$ $73.7$ $99.3$ $40.9$
Google T5-Base $79.7$ $75.9$ $99.6$ $53.1$
Several popular studies specifically address defenses against backdoor attacks in NLP. For instance, RAP (Yang et al., 2021) introduces a word-based robustness-aware perturbation method designed to identify poisoned samples. And ONION (Qi et al., 2020) attempts to remove trigger words by empirically assessing sentence perplexities. However, they cannot handle our TrojFSP using invisible syntactic triggers.
We propose a potential defense technique against TrojFSP that minimizes its ASR by selectively pruning unimportant prompt tokens, assuming the defender knows the potential presence of a backdoored prompt. However, the token’s importance may vary with different input samples. Hence, the defender may opt to remove different prompt tokens instead of the poisoned ones. Therefore, even after token pruning, TrojFSP can still achieve an ASR of over $40\%$, as demonstrated in Table 9. There is a need for a more efficient and accurate defense method.
@@ 6 Conclusion
In this paper, we propose a new backdoor attack, TrojFSP, on few-shot prompt-tuning with TC-Shrink, selective token poisoning and Trojan-Trigger Attention. We also discuss the potential defense techniques in this paper. Compared to prior prompt-based backdoor attacks, TrojFSP improves the ASR by $9\%\sim 48\%$ and the CDA by $4\%\sim 9\%$ across various PLMs and a wide range of downstream tasks.
7 Limitation
The limitations of our paper are as follows: (i) Dataset and Tasks. Our TrojFSP evaluates few-shot prompt attacks on popular benchmark datasets and models, including SST-2, SST-5, MR, Twitter, and LingSpam datasets; RoBERTa-Large, GoogleT5-Base, and GPT-J models. However, the paper primarily focuses on classification tasks, potentially constraining the generalizability of our findings to a broader range of NLP tasks such as generation Chen et al. (2023). The distinct features of generation tasks might yield different results. (ii) Attacking Conditions. While attackers can easily obtain and implement attacks on models from shared platforms and open-sourced models, the landscape of black-box attacks, particularly those involving APIs, remains unclear. Investigating these scenarios is an avenue for future research.
8 Ethics Statement
In this paper, we present research on the potential threat posed by few-shot prompt backdoor attacks, showcasing their effectiveness and stealthiness. Our primary objective is to raise awareness among NLP practitioners about the risks associated with untrusted prompts and to stimulate further research aimed at mitigating the threat of backdoor attacks in NLP systems.
While we acknowledge that the malicious use of the proposed attack method could raise ethical concerns, including potential security risks and issues related to trust in NLP systems, it is important to emphasize several factors that limit the real-world harm of our proposed approach. These limitations include the usage of poisoned prompt and trigger types.
Additionally, we contribute to ethical considerations by proposing a defense method against the described attack. This defense mechanism is intended to assist in minimizing the potential harm and enhancing the robustness of NLP systems against backdoor attacks.